log10(6x-3)=4
solve
I have no clue

Found 2 solutions by ewatrrr, nyc_function:
Answer by ewatrrr(24785)


Hi,

log10(6x-3)=4
10^4 = 6x-3
1667.167 = x

Answer by nyc_function(2741)

log_10 (6x - 3) = 4

(1) Rewrite the equation in exponential form.

10 is raised to the 4th power and (6x - 3) moves to the right side of the equation.

10^4 = 6x - 3

(2) We now solve equation for x.


10,000 = 6x - 3

10,000 + 3 = 6x

10003 = 6x

(3) We now divide 10003 by 6 to find x.

10003/6 = x

1667.17 = x